OpenStudy (anonymous):

(n^2+3n+2)/(2)÷(n+2)/(2(n+2))

OpenStudy (anonymous):

factor that trinomial in the first fraction

OpenStudy (anonymous):

what would the answer be, im stuck :)

OpenStudy (anonymous):

thank guys i got it :) xoxoxxx

OpenStudy (anonymous):

Is this what you wanted simplified \[ \frac{\frac{n^2+3 n+2}{2}}{{\frac{ (n+2)}{2 (n+2)}}} \]

OpenStudy (anonymous):

If this is the case is answer is n^2+3n+2